52.02Attendance is up at the local minor league stadium this year. Last year there was an average of 2,918 fans per game. This year the average has been 4,436. What percent increase has occurred? Round your answer to the nearest hundredth of a percent.
Answer by nyc_function(2741)

(increase amount)/(original amount) = p/100, where p = percent increase
(4436 - 2918)/2918 = p/100
1518/2918 = p/100
2918p = 1518 * 100
2918p = 151800
p = 151800/2918
p = 52.02%